# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v1

@@ 5 Method
[…]
  1. (i)
    Initialize the ViT with $W_{D_{P}}$ from a large pre-training dataset (eg: DinoV2 or MAE weights).
  2. (ii)
[…]
  3. (iii)
    For all other blocks ${\mathcal{L}}\setminus{\mathcal{U}}$, freeze all parameters and use LoRA with rank $r$ on the $Q$ and $V$ weights of the attention layer. Also unfreeze the normalization layers of these blocks.
  4. (iv)
    Train the resulting model on unlabeled images ${\mathbf{x}}\sim p_{D_{F}}(\cdot)$ of the new domain $D_{F}$, with the same unsupervised objective as what was used for $W_{D_{P}}$ (eg: DinoV2 or MAE training objective).
In terms of notation, D-ExPLoRA-[L]-$r32$ would then denote a ViT initialized with DinoV2 weights (as opposed to M, which is MAE weights), where ${\mathcal{U}}=\{\texttt{L}\}$, and LoRA rank 32 is used on the $Q,V$ matrices of every attention layer in ${\mathcal{L}}\setminus{\mathcal{U}}$. In this way, $\Delta_{D_{F}}$ requires 5-10% of the original parameters of the ViT, and we show it learns unsupervised representations for $\Delta_{D_{F}}$ that can match or even outperform learning $W_{D_{F}}$ directly from scratch, through full pre-training.
@@
For our experiments, we use the DinoV2 ViT-L model as $W_{D_{P}}$, without registers [47]. We unfreeze the 24th block of the ViT, and use LoRA only on the query $Q$ and value $V$ matrices of each attention layer in all other blocks. We also unfreeze the normalization layers (which require very few parameters) throughout the network. We train each of the Dino, iBOT, and Koleo linear heads fully, without any frozen parameters.
@@
We initialize both the MAE encoder and decoder with pre-trained weights. We only unfreeze the last block of the ViT encoder, and tune the $Q,V$ matrices of each attention layer in all other blocks (including all blocks of the decoder) with LoRA.
For the multi-spectral MAE introduced in [22] we need to unfreeze the patch embedding layers for each group of channels (as these cannot be initialized from $W_{D_{P}}$ since $W_{D_{P}}$ only considers 3 channel RGB inputs). We then find that when using ExPLoRA, unfreezing blocks 1 and L is necessary to successfully achieve domain transfer.
Fine-tuning
While fine-tuning DinoV2, we discard the linear heads used for the pre-training loss components (section 3), and load all other model weights. Similarly, for MAE, we discard the decoder weights. We then initialize a linear head for classification, or decoder for segmentation, either of which is fully trainable. Freezing all ViT encoder weights, we then use LoRA-$r8$ on the $Q,V$ matrices of every attention layer. We find that the drop-path augmentation [48] is especially useful for fine-tuning, and use a value of 0.2 for the ViT-L models.
@@ 6 Experiments
Our experimental results first consist of a case study on satellite imagery (section 6.1), including an ablation study in section 6.1.2. We then further evaluate on downstream tasks in sections 6.1.3, 6.1.4 and 6.2. Training details including hyperparameter and compute configurations are mentioned in appendix B. Our results achieve a new SoTA top 1 accuracy of 79.1% ($\uparrow$1.3%) on the competitive fMoW-RGB benchmark, outperforming fully pre-trained and fine-tuned models while using 6% of the ViT encoder parameters. We also achieve a $\uparrow$7.3improvement in linear probing accuracy on the same dataset. Across other satellite datasets, we match fully-pretrained prior state-of-the-art methods, and demonstrate competitive performance on WiLDS benchmark datasets as well.
@@ 6.1 Case Study: Satellite Imagery
We begin by examining satellite images largely because of the proliferation in works developing foundational models for satellite imagery via pre-training from scratch[22, 23, 24, 25] and since they represent a significant domain shift from natural images. We thus structure our study in a few parts: (i) Evaluating ExPLoRA on RGB satellite images, both for linear probing and PEFT fine-tuning (ii) Demonstrating ExPLoRA ability to handle a domain shift from natural images to multi-spectral and temporal satellite inputs, a more difficult translation. (iii) Evaluating models pre-trained with ExPLoRA on downstream satellite datasets.
@@ 6.1.1 RGB Satellite Images
We first consider the functional map of the world (fMoW) dataset of high-resolution satellite images, each paired with one of 62 classification labels [49]. This dataset is used as a benchmark for multiple satellite-image foundation models [22, 23, 24].
Model Backbone PEFT
@@
MAE [7] ViT-L Full - 303.3M 76.91
SatMAE [22] ViT-L LoRA-r8 303.3M 0.8M 76.10
MAE [7] ViT-L LoRA-r8 - 0.8M 76.21
MAE [7] ViT-L BOFT-b2m8 - 0.9M 72.40
DinoV2 [6] ViT-L LoRA-r8 - 0.8M 78.08
M-ExPLoRA-[L]-$r64$ ViT-L LoRA-r8 18.7M 0.8M 76.55
D-ExPLoRA-[L]-$r64$ ViT-L LoRA-r8 18.7M 0.8M 79.15
Table 1: Results on the fMoW-RGB validation dataset. The “Pre-train #Params" and “Fine-tune #Params" refer to the trainable parameters of the ViT encoder required on the new domain, i.e. satellite images.
We compare our results in table 1 against both prior fully pre-trained SoTA foundation models as well as PEFT techniques applied on ViTs pre-trained with MAE and/or DinoV2 weights. Our results demonstrate that D-ExPLoRA-$\tt{[L]}$-$r64$ is SoTA in terms of fMoW-RGB average accuracy at 79.15%. ExPLoRA outperforms techniques that require fully pre-training ViTs on fMoW while using 6% of the original ViT encoder parameters. We also note that the original LoRA configuration outperforms other PEFT techniques when paired with the drop-path regularization technique. For example, we find that BOFT does not pair well with drop-path, instead performing most effectively with a custom multiplicative dropout technique. We include the result with the best hyperparameter configuration for each row in table 1.
Method Backbone Top 1 Acc.
GASSL [23] ResNet50 68.32
@@
D-ExPLoRA-[L]-$r64$ ViT-L 75.77
D-ExPLoRA-[L]-$r64$$\dagger$ ViT-L 76.53
Table 2: Linear-probing on fMoW-RGB. The first four rows fully pre-train on the dataset. $\dagger$ denotes concatenating features from the last 4 ViT blocks.
Using our strongest performing variant (i.e. ExPLoRA with DinoV2), we investigate linear-probing performance on fMoW-RGB compared with prior SoTA methods in table 2. Linear-probing represents freezing the backbone and then training a linear head on the features extracted from the frozen backbone, serving as a desirable metric of the quality of extracted embeddings. Our results demonstrate an improvement of over $\uparrow$7.3in top 1 average accuracy over prior SoTA methods, demonstrating that ExPLoRA learns robust unsupervised representations for its target domain without requiring expensive from-scratch pre-training. Importantly, ExPLoRA outperforms domain-specific prior SoTA solutions (rows 1-4), as well as DinoV2, which suggests successful transfer learning on the target domain by leveraging knowledge from pre-training on natural images.
6.1.2 Ablation study
Figure 3: Lin. probe accuracy vs. number of training iterations, demonstrating ExPLoRA’s efficiency even with few pre-training updates.
@@ 6.1.2 Ablation study
[1] 32 [Q,V] 15.7M 73.39
[L] 32 [Q,V] 15.6M 75.14
[L] 8 [Q,V] 13.4M 75.23
[L] 32 [Q,V] 15.7M 75.44
[L] 64 [Q,V] 18.7M 76.53
Table 3: Ablation study using DinoV2-ExPLoRA, measuring linear-probing accuracy on fMoW-RGB. If the LoRA rank is $>0$, LoRA is only used on the frozen ViT blocks. All results are obtained by using concatenated features from the last 4 ViT blocks.
We perform an ablation study (table 3) on linear-probing performance for fMoW-RGB to determine whether our proposed configuration performs optimally. A natural question is whether the improvement in performance stems primarily from unfreezing blocks, or from LoRA-tuning the rest of the ViT. We investigate this by unfreezing blocks L,L-1 in the row 1 (with no LoRA), and comparing that with ExPLoRA-L-$r8$ in row 6. As seen, unfreezing an extra block consumes almost double the number of parameters, but fails to yield the same improvement in performance. Thus, simply increasing the number of unfrozen blocks will likely improve performance, but will not do so as effectively as ExPLoRA, and will also significantly and sharply decrease the parameter-efficiency. As seen in fig. 3, ExPLoRA also enables faster convergence (blue) with fewer training iterations compared to unfreezing more blocks (red).
Next, we investigate whether high LoRA ranks used on all ViT layers (i.e. all attention and MLP matrices, not just $Q,V$) is beneficial. Surprisingly, this significantly harms learning (row 3). In fact, it is much less effective than using just LoRA-$r256$ on the $Q,V$ matrices of all ${\mathcal{L}}$ blocks (row 2). However, both rows 2 and 3 are much less parameter-efficient than ExPLoRA (rows 4-6).
@@
Lastly, we investigate the impact of LoRA rank on ExPLoRA. Changing the rank from 8 to 32 has a small improvement, but changing from 32 to 64 brings about a much larger improvement, with only a relatively small increase in trainable parameters. This demonstrates that higher ranks are necessary during pre-training for effective learning on the new domain. One hypothesis for the effectiveness of pairing unfreezing blocks with LoRA tuning is that the low-rank updates to the ViT backbone “nudge" the sequence of embedded visual tokens from $D_{P}$ to those representing $D_{F}$, which then enables the unfrozen ViT block to effectively compress data from the new domain.
@@ 6.1.3 Multi-Spectral Satellite Images
Next, we consider the fMoW-Sentinel dataset, a large dataset of Sentinel-2 images used in [22]. Each image consists of 13 spectral bands and is paired with one of 62 classes.
With fMoW-Sentinel, we aim to assess the feasibility of domain transfer from natural images to multi-spectral, low-resolution satellite images. This presents a significant challenge compared to fMoW-RGB, as none of the natural image datasets in $D_{P}$ include sensor information beyond visible light RGB bands. We utilize the group-channel ViT-L SatMAE model introduced in [22], initializing it with MAE self-supervised weights from the natural image domain. Since the patch embedding layers differ from those of MAE, we unfreeze and train them from scratch during ExPLoRA pre-training, resulting in minimal overhead to the parameter count.
@@
From table 4, we observe the challenge of domain transfer from natural images to multi-spectral satellite images, as discussed in section 5. Even fully fine-tuning from MAE weights results in nearly a 10% drop in accuracy (row 2). LoRA tuning solely from MAE weights (row 4) performs even worse. Ablating by initializing with MAE weights and unfreezing only 4 transformer blocks during pre-training (row 6) is insufficient to bridge the domain gap. Notably, with ExPLoRA, unfreezing the first and last transformer blocks yields surprisingly good results, surpassing even fully pre-training from scratch (when using LoRA for fine-tuning). This underscores ExPLoRA’s ability to bridge wide domain gaps while utilizing only a fraction (in this case, around a tenth) of the original parameters.
@@ 6.1.4 Additional Satellite Datasets
Also sourced from fMoW-RGB [49], each input is a sequence of up to 3 images of the same location, distributed temporally, and paired with one of 62 classes. Since the inputs are now temporal sequences, we initialize the temporal MAE architecture from [22] with MAE weights, and pre-train on the dataset’s training images (without labels) with ${\mathcal{U}}=\tt[L]$ and LoRA rank 32. Our LoRA-tuned model then outperforms the domain-specific SatMAE for PEFT (table 8), demonstrating successful transfer learning at a fraction of the pre-training parameters used by temporal SatMAE, which was fully pre-trained and fully fine-tuned on this dataset.
EuroSAT
The dataset contains 27,000 13-band satellite images of 10 classes [50], sourced from Sentinel-2. For ExPLoRA, we don’t pre-train on the training set of this dataset, and instead use LoRA fine-tuning starting with the pre-trained weights learned in row 8 of table 4. We demonstrate improved performance over DinoV2, and match the performance achieved by the domain-specific SatMAE which was fully pre-trained on fMoW-Sentinel, and fully fine-tuned on EuroSAT (table 8). This demonstrates the successful use of our extended pre-trained model on further downstream datasets.
@@
This dataset contains high resolution satellite images, each paired with a segmentation mask for buildings [51]. The training and test sets consist of 5000 and 1940 images, respectively. For ExPLoRA, we pre-train on the training set. A significant portion of the data consists of images with extensive black regions, indicating areas without meaningful visual information. Considering this limitation and the small dataset size, it is not clear whether additional pretraining is effective. We find that, despite this, ExPLoRA remains on par with the LoRA tuned DinoV2 model and remains competitive with the fully pre-trained and fully fine-tuned domain-specific models (table 8).
@@
The benchmark RESISC-45 [52] dataset consists of 31,500 satellite images of varying resolution (0.2m-30m GSD), with 45 classes. The data is split into 25,200 training and 6,300 validation images, as per Reed et al. [24]. In table 8, our D-ExPLoRA pre-trained on only high-resolution fMoW-RGB images achieves SoTA results of 97.16% on multi-resolution RESISC-45 images, with just linear-probing. Since we use the same pre-trained model as in the last row of table 1, we demonstrate successful transfer learning from ExPLoRA pre-training, without requiring any additional modifications for scale-aware representation learning [24].
Method PEFT Top 1 Acc.
@@
Method PEFT Top 1 Acc.
SeCo [53] Full 93.14
SatMAE [22] Full 98.98
SatMAE [22] LoRA-r8 98.73
DinoV2 [6] BOFT-b8m2 96.60
M-ExPLoRA-[1,L]-$r32$ LoRA-r8 98.54
Table 6: EuroSAT validation set results
Method PEFT mIoU
GASSL [23] Full 78.51
SatMAE [22] Full 78.07
ScaleMAE [24] Full 78.90
DinoV2 [6] Lin. Probe 76.21
DinoV2 [6] LoRA-r8 76.69
D-ExPLoRA-[L]-$r64$ LoRA-r8 76.69
Table 7: SpaceNet validation set results
Method PEFT Top 1 Acc.
SatMAE [22] Full 94.80
ScaleMAE [24] Full 95.70
GASSL [23] Lin. Probe 89.94
SatMAE [22] Lin. Probe 88.30
ScaleMAE [24] Lin. Probe 89.60
DinoV2 [6] Lin. Probe 96.34
D-ExPLoRA-[L]-$r64$ Lin. Probe 97.16
Table 8: Resisc-45 validation set results

### v2

@@ 5 Method
To learn $\Delta_{D_{T}}$, we propose ExPLoRA (i.e. Extended Pre-training with LoRA), a method that efficiently adapts pre-trained ViTs for new target domains $D_{T}$, described in algorithm 1.
In terms of notation, D-[L]-$r64$ refers to a ViT initialized with DinoV2 weights (denoted by D), where ${\mathcal{U}}=\{\texttt{L}\}$, and LoRA rank 64 is applied to the $Q,V$ matrices of attention layers in ${\mathcal{L}}\setminus{\mathcal{U}}$. Thus, $\Delta_{D_{T}}$ comprises of all weights in ${\mathcal{U}}$, LoRA matrices in ${\mathcal{L}}\setminus{\mathcal{U}}$, and normalization layers. For ${\mathcal{U}}=\{\texttt{L}\}$, $\Delta_{D_{T}}$ consists of only 5% of the original ViT parameters. As we show in section 6, our extended pre-training approach can match or even outperform full pre-training on new domains from scratch.
@@
We initialize a ViT-L with $W_{D_{S}}$ from the DinoV2 ViT-L encoder, without registers (Darcet et al., 2023). Since the DinoV2 pre-trained checkpoints don’t contain the Dino or iBOT linear heads, we initialize a shared Dino-iBOT linear head from scratch. This shared head is fully trained during extended pre-training, adding only a minimal number of trainable parameters.
@@
We initialize a ViT-L with $W_{D_{S}}$ from the MAE ViT-L encoder. Since MAE provides the pre-trained decoder, we use these weights to initialize our MAE decoder He et al. (2022). During extended pre-training, in addition to the ExPLoRA recipe in algorithm 1, we apply LoRA with rank $r^{\prime}$ on the $Q,V$ matrices of each attention layer in the frozen decoder. Note that the LoRA rank $r^{\prime}$ may differ from the LoRA rank $r$ used in the ViT encoder (section A.4). All other decoder weights, apart from the layer-normalization layers, are kept frozen. No block is fully unfrozen in the MAE decoder, as it will be discarded after extended pre-training. This helps to minimize the number of additional parameters trained in the decoder.
ExPLoRA for Multi-Spectral Inputs
For the multi-spectral ViT introduced by SatMAE we need to additionally unfreeze the positional encoding and the patch embedding weights for each group of channels. These cannot be initialized from $W_{D_{S}}$, as $W_{D_{S}}$ is trained on RGB inputs, whereas multi-spectral inputs can have more or different channels. As part of $\Delta_{D_{T}}$ in algorithm 1, the positional encodings and patch embeddings for multi-spectral data are adapted during extended pre-training. Aside from this, the approach remains unchanged from that of DinoV2 or MAE described earlier.
Storage Considerations
After running ExPLoRA, we receive a new unsupervised model $W_{D_{T}}^{\ast}=W_{D_{S}}+\Delta_{D_{T}}$ for the target domains $D_{T}$. Any components that are not part of the ViT encoder (eg: the Dino linear head or the MAE decoder) are discarded. Post-ExPLoRA, only $\Delta_{T}$, consisting of 1-2 unfrozen ViT blocks, LoRA matrices, and layer-normalization weights, are stored for each $D_{T}$. Like LoRA and other PEFT methods, ExPLoRA significantly reduces additional storage requirements compared to fully training $W_{D_{T}}$ from scratch.
Fine-Tuning post-ExPLoRA
After extended pre-training with ExPLoRA, the output weights $W_{D_{T}}^{\ast}$ behave as any fully pre-trained model $W_{D_{T}}$. We can now use the ViT for feature extraction, PEFT, or fine-tuning as desired. For instance, we could initialize a linear head for classification or a decoder for segmentation, either of which is fully trainable. We can then freeze all ViT weights and apply LoRA-$r8$ on the $Q,V$ matrices of the attention layers (or use another PEFT method). Lastly, we use supervised fine-tuning on each labeled dataset ${\mathcal{Y}}_{d_{T}}$ to train the unfrozen parameters $\Delta_{d_{T}}$. This yields our final model $W_{d_{T}}$ (eq. 4), which can be used for classification, segmentation, detection etc.
@@ 6 Experiments
Our experimental results consist of a case study on satellite imagery (section 6.1), with an ablation study in section 6.1.2 and analysis in section 6.3. We evaluate on multiple downstream tasks in sections 6.1.3, 6.1.4 and 6.2. Additional experiments and ablations are provided in appendix A and training hyperparameter and compute configurations are mentioned in appendix B. Our results achieve a new SoTA top 1 accuracy of 79.2% ($\uparrow$1.4%) on the competitive fMoW-RGB benchmark, outperforming fully pre-trained and fine-tuned models while using 6% of the ViT encoder parameters. We also achieve a $\uparrow$8.2improvement in linear probing accuracy on the same dataset. Across other satellite datasets, we match fully-pretrained prior state-of-the-art methods, and demonstrate competitive performance on WiLDS benchmark datasets as well.
@@ 6.1 Case Study: Satellite Imagery
We examine satellite images given their importance towards multiple societal applications (Broader Impact) and since they represent a significant domain shift from natural images. There is a large and growing body of research on developing foundational models for satellite imagery from scratch Ayush et al. (2021); Cong et al. (2022); Reed et al. (2023); Tang et al. (2024), thus presenting a good benchmark for ExPLoRA.
@@ 6.1.1 RGB Satellite Images
We first consider the functional map of the world (fMoW) dataset of high-resolution satellite images, each paired with one of 62 classification labels (Christie et al., 2018). fMoW is used as a benchmark for many satellite-image foundation models (Cong et al., 2022; Reed et al., 2023).
Model Arch. PEFT
@@
SatMAE [15] ViT-L LoRA-$r8$ [28] 303.3M 0.8M 76.10
MAE [26] ViT-L LoRA-$r8$ [28] - 0.8M 76.21
MAE [26] ViT-L DVPT-10 [30] - 0.4M 72.35
MAE [26] ViT-L GVPT-100 [60] - 0.4M 70.86
MAE [26] ViT-L SA2VP [47] - 1.1M 73.55
MAE [26] ViT-L AdaLoRA-$r8$ [62] - 1.2M 75.25
M-[L]-$r64$ ViT-L LoRA-$r8$ [28] 18.7M 0.8M 76.55
DinoV2 [45] ViT-L LoRA-$r8$ [28] - 0.8M 78.08
DinoV2 [45] ViT-L BOFT-$b2m8$ [40] - 0.9M 72.40
DinoV2 [45] ViT-L VPT-100 [30] - 0.4M 77.29
DinoV2 [45] ViT-L GVPT-100 [60] - 0.4M 76.22
DinoV2 [45] ViT-L AdaLoRA-$r8$ [62] - 1.2M 78.87
DinoV2 [45] ViT-L SA2VP [47] - 1.1M 77.53
D-[L]-$r64$ ViT-L SA2VP [47] 18.7M 1.1M 78.51
D-[L]-$r64$ ViT-L LoRA-$r8$ [28] 18.7M 0.8M 79.15
Table 1: Results on the fMoW-RGB validation dataset. The “Pre-train #Params" and “Fine-tune #Params" refer to the trainable parameters of the ViT encoder required on the new domain, i.e. satellite images. M-[L]-$r64$ and D-[L]-$r64$ refer to ExPLoRA models initialized with MAE and DinoV2 weights, respectively (section 5).
We compare our results in table 1 against both prior fully pre-trained SoTA foundation models as well as PEFT techniques applied on ViTs pre-trained with MAE and/or DinoV2 weights. Our results demonstrate that D-ExPLoRA-$\tt{[L]}$-$r64$ is SoTA in terms of fMoW-RGB average accuracy at 79.15%. ExPLoRA outperforms techniques that require fully pre-training ViTs on fMoW while using 6% of the original ViT encoder parameters.
ExPLoRA-initializations with LoRA fine-tuning outperform other unsupervised initializations paired with PEFT techniques by 1-3%, including SoTA matrix-adaptation methods like AdaLoRA (Zhang et al., 2023b), BOFT (Liu et al., 2023), and VPT-based approaches such as GVPT (Yoo et al., 2023) and SA2VP (Pei et al., 2024). Additionally, applying SA2VP to ExPLoRA-initialized ViTs improves performance over a DinoV2 initialization by  1%, showcasing ExPLoRA’s compatibility with other PEFT methods and its versatility as an initialization for new domains.
Using our strongest performing variant (i.e. ExPLoRA with DinoV2), we investigate linear-probing performance on fMoW-RGB compared with prior SoTA methods in table 3. Linear-probing represents freezing the backbone and then training a linear head on the features extracted from the frozen backbone, serving as a desirable metric of the quality of extracted embeddings. Our results demonstrate an improvement of over $\uparrow$8.2in top 1 average accuracy over prior SoTA methods, demonstrating that ExPLoRA learns robust unsupervised representations for its target domain without requiring expensive from-scratch pre-training. Importantly, ExPLoRA outperforms domain-specific prior SoTA solutions (rows 1-4), as well as DinoV2, which suggests successful transfer learning on the target domain by leveraging knowledge from pre-training on natural images.
6.1.2 Ablation study
Method Arch.
Top 1
Acc.
GASSL [2] ResNet 68.32
@@ 6.1.2 Ablation study
D-[L]-$r64$ ViT-L 76.86
D-[L]-$r64$$\dagger$ ViT-L 77.48
Table 2: Linear-probing on fMoW-RGB. The first four rows fully pre-train on the dataset. $\dagger$ denotes concatenating features from the last 4 ViT blocks. All other rows use the features of the last ViT block.
@@
[1] 64 [Q,V] 18.7M 75.97
[9] 64 [Q,V] 18.7M 75.45
[L-1] 64 [Q,V] 18.7M 77.40
[L] 0 VPT-100 12.8M 70.14
[L] 64 [Q,V] 18.6M 76.78
[L] 8 [Q,V] 13.4M 76.31
[L] 32 [Q,V] 15.7M 76.40
[L] 64 [Q,V] 18.7M 77.48
Table 3: Ablation study using DinoV2-ExPLoRA, measuring linear-probing accuracy on fMoW-RGB. All results are obtained by using concatenated features from the last 4 ViT blocks.
@@
Next, we investigate whether high LoRA ranks used on all ViT layers (i.e. all attention and MLP matrices, not just $Q,V$) is beneficial. Surprisingly, this significantly harms learning (row 4). In fact, it is much less effective than using just LoRA-$r256$ on the $Q,V$ matrices of all ${\mathcal{L}}$ blocks (row 3). However, both rows 3 and 4 are much less parameter-efficient than ExPLoRA (rows 5-7, 10-12).
[…]
Lastly, we investigate the impact of LoRA rank on ExPLoRA. Changing the rank from 8 to 32 has a small improvement ($\uparrow 0.09\%$), but changing from 32 to 64 brings about a much larger improvement ($\uparrow 1.08\%$), with only a relatively small increase in trainable parameters. This demonstrates that higher ranks are necessary during pre-training for effective learning on the new domain. Further ablations on data efficiency (A.3), MAE decoder rank (A.4), and ViT backbone size (A.5) are in appendix A.
@@ 6.1.3 Multi-Spectral Satellite Images
Next, we consider the fMoW-Sentinel dataset, a large dataset of Sentinel-2 images used in Cong et al. (2022). Each image consists of 13 spectral bands and is paired with one of 62 classes.
With fMoW-Sentinel, we evaluate the challenge of transferring from natural images to multi-spectral, low-resolution satellite images– a much harder task than fMoW-RGB due to the absence of sensor data beyond RGB bands in $D_{S}$. We use the SatMAE group-channel ViT-L model from Cong et al. (2022), initialized with MAE weights from natural images. Since the patch embedding layers differ, we unfreeze and train them from scratch during algorithm 1, adding minimal parameter overhead.
@@
Table 4 highlights the difficulty of this domain transfer. Fully fine-tuning from MAE weights leads to a nearly 10% drop in accuracy (row 2), and LoRA tuning from MAE alone performs even worse (row 4). Simply unfreezing four transformer blocks during pre-training (row 6) fails to close the gap. However, ExPLoRA, with ${\mathcal{U}}=\{\tt{1,L}\}$, yields excellent results, even outperforming full pre-training from scratch for LoRA fine-tuning (row 5 vs. last row). This demonstrates ExPLoRA’s effectiveness in bridging large domain gaps while using just a fraction (here, <10%) of the parameters.
@@ 6.1.4 Additional Satellite Datasets
We perform extensive experiments on downstream satellite datasets, with further results in A.1.
@@
Each input is a sequence of up to 3 fMoW-RGB (Christie et al., 2018) images of the same location, distributed temporally, and paired with one of 62 classes. Since the inputs are now temporal sequences, we initialize the temporal MAE architecture from Cong et al. (2022) with MAE weights, and pre-train on ${\mathcal{X}}_{D_{T}}$ with ${\mathcal{U}}=\tt[L]$ and LoRA rank 32. ExPLoRA then outperforms temporal SatMAE for PEFT (table 6), demonstrating successful transfer learning at a fraction of the pre-training parameters.
@@
This dataset contains high resolution satellite images, each paired with a segmentation mask for buildings (Van Etten et al., 2018). The training and test sets consist of 5000 and 1940 images, respectively. For ExPLoRA, we pre-train on the training set. However, many images in the dataset contain extensive blacked-out regions, indicating limits of the visible region. Considering this limitation and the small dataset size, it is not clear whether additional pre-training is effective. We find that, despite this, ExPLoRA remains on par with the LoRA-tuned DinoV2 model and remains competitive with the fully pre-trained and fully fine-tuned domain-specific models (table 6).
@@
The RESISC-45 (Cheng et al., 2017) benchmark dataset consists of 31,500 satellite images of varying resolution (0.2m-30m GSD), with 45 classes. The data is split into 25,200 training and 6,300 validation images, as per Reed et al. (2023). In table 6, our D-ExPLoRA pre-trained on only high-resolution fMoW-RGB images achieves SoTA results of 97.32% on multi-resolution RESISC-45 images, with just linear-probing. Since we use the same pre-trained model as in the last row of table 1, we demonstrate successful transfer learning from ExPLoRA pre-training, without requiring any additional modifications for scale-aware representation learning (Reed et al., 2023).
Method PEFT
Top 1 Acc.
@@
Method PEFT
SpaceNet
mIoU
Resisc45
Top 1 Acc.
SatMAE [15] Full 78.07 94.80
ScaleMAE [51] Full 78.90 95.70
DinoV2 [45] LoRA-r8 76.69 97.60
D-[L]-$r64$ LoRA-r8 76.69 97.65
SatMAE [15] Lin. Probe 50.89 88.30
ScaleMAE [51] Lin. Probe 47.17 89.60
DinoV2 [45] Lin. Probe 76.21 96.34
D-[L]-$r64$ Lin. Probe 76.34 97.32
Table 6: SpaceNet and Resisc-45 validation set results